Machinery at a factory originally costing $50,000 depreciates $1800 the first year, $1750 the second year, $1700 the third year, etc.
What is the amount of depreciation during the fifth year?
What is th value of the machinery at the end of the fifth year?

Answer by Jstrasner(112)

Hey,
so for this one we need to find the sum of the depreciations of the money over the course of 5 years:
The first year, the machinery depreciates 1800, the next year 1750 and the next year 1700. First we see the pattern that every year the depreciation goes down $50. So now we need to find out the depreciation for the 4th and 5th year. The 3rd year, the depreciation is $1700. If the depreciation goes down $50. The depreciation for the 4th year is 1700-50 = 1650. The depreciation for the 5th year is 1650-50 = 1600.
Now we know the amount of depreciation during the 5th year: $1600.
The value of the machinery at the end of the fifth year is found by subtracting the total amount of depreciation from the original cost:
50,000 - sum of depreciation = value
We now need to find the sum of the 5 years:
1800 + 1750 + 1700 + 1650 + 1600 = 8500
So the value at the end of the fifth year is:
$50,000 - $8,500 = $41,500
I hope this helps!